1. Crane Marine Products began the year with 10 units of marine floats at a cost of $15.00 each. During the year, it made the following purchases: May 5, 30 units at $22.00; July 16, 15 units at $27.00; and December 7, 20 units at $33.00. Assuming there are 25 units on hand at the end of the period, determine the cost of goods sold under (a) FIFO, (b) LIFO, and (c) average-cost. Crane uses the periodic approach.

Cost of Goods Sold

FIFO

$enter cost of goods sold under First In First Out in dollars

LIFO

$enter cost of goods sold under Last In First Out in dollars

Average-cost

$enter cost of goods sold under Average-cost in dollars

2. Suppose at December 31 of a recent year, the following information (in thousands) was available for sunglasses manufacturer Oakley, Inc.: ending inventory $170,500, beginning inventory $132,000, cost of goods sold $378,125, and sales revenue $836,000.

Calculate the inventory turnover and days in inventory for Oakley, Inc. (Round inventory turnover to 2 decimal places, e.g. 15.25 and days in inventory to 0 decimal places, e.g. 15. Use 365 days for calculation.)

Inventory turnover

enter inventory turnover in times rounded to 2 decimal places

times

Days in inventory

enter a number of days in inventory rounded to 0 decimal places

days

3. Use a tabular summary to record the following transactions for Marin Co. (Omit recording cost of goods sold.)

(a) On July 1, Marin Co. sold merchandise on account to Stacey Inc. for $20,820.
(b) On July 8, Stacey Inc. returned $2,520 of July 1 merchandise to Marin Co.
(c) On July 11, Stacey Inc. paid for the merchandise.


(Enter negative amounts using either a negative sign preceding the number e.g. -45 or parentheses e.g. (45).)

Assets

=

Liabilities

+

Stockholders' Equity

Retained Earnings

Cash

+

Accts. Rec.

= +

Common Stock

+

Rev.

-

Exp.

(a) July 1 $enter a dollar amount $enter a dollar amount $enter a dollar amount $enter a dollar amount $enter a dollar amount select an account title                                                          Sales returns & allowancesCost of goods soldAccounts receivableBad debt expenseSales revenue
(b) 8 enter a dollar amount enter a dollar amount enter a dollar amount enter a dollar amount enter a dollar amount select an account title                                                          Bad debt expenseCost of goods soldSales returns & allowancesAccounts receivableSales revenue
(c) 11 enter a dollar amount enter a dollar amount enter a dollar amount enter a dollar amount enter a dollar amount select an account title                                                          Bad debt expenseCost of goods soldSales revenueAccounts receivableSales returns & allowances
Answer #1

1.
Cost of goods sold:

FIFO = $1,080
LIFO = $1,395
Average Cost = $1,250

Units Per unit Total Cost 10 $15.00 $150.00 FIFO - Periodic Units available for sale Opening inventory Purchases May 5 July 1


Average cost periodic method

Total units of goods available for sale = 75 units

Total cost of goods available for sale = $1,875.00

Average cost = $1,875.00/ 75 = $25 per unit

Ending inventory = 25 units

Units sold = 75 - 25 = 50 units

Cost of goods sold = Units sold × Average cost per unit = 50 × $25 = $1,250

2.
Inventory turnover = 2.5 times
Days in inventory = 146 days

Inventory turnover = Cost of goods sold / Average inventory

Cost of goods sold = $378,125
Average inventory = (Beginning inventory + Ending inventory)/ 2
= ($132,000 + $170,500) / 2 = $151,250

Inventory turnover = $378,125/ $151,250 = 2.5

Days in inventory = 365 / Inventory turnover = 365 / 2.5 = 146 days

3.

Assets Liabilities + Cash + Accts. Rec. + + + Stockholders Equity Common + Retained Earnings Stock Rev. Exp. . $20,820 $2,52

Goods sold on account is debit to accounts receivable and credit to sales revenue.

The merchandise returned is debited as sales returns and allowances and credited as accounts receivable.

For Cash received, debit cash and credit accounts receivable.